### v1

# Attention-Based Multi-Scale Temporal Fusion Network for Uncertain-Mode Fault Diagnosis in Multimode Processes

Abstract
Fault diagnosis in multimode processes plays a critical role in ensuring the safe operation of industrial systems across multiple modes. It faces a great challenge yet to be addressed – that is, the significant distributional differences among monitoring data from multiple modes make it difficult for the models to extract shared feature representations related to system health conditions. In response to this problem, this paper introduces a novel method called attention-based multi-scale temporal fusion network. The multi-scale depthwise convolution and gated recurrent unit are employed to extract multi-scale contextual local features and long-short-term features. A temporal attention mechanism is designed to focus on critical time points with higher cross-mode shared information, thereby enhancing the accuracy of fault diagnosis. The proposed model is applied to Tennessee Eastman process dataset and three-phase flow facility dataset. The experiments demonstrate that the proposed model achieves superior diagnostic performance and maintains a small model size.
keywords:
Fault diagnosis , Multimode process , Deep learning
journal: Elsevier
[…]
1 Introduction
With the rapid development of automation, sensor and computer technologies, the scale and complexity of industrial systems have grown significantly. The complex industrial systems pose significant challenges in terms of safety and reliability engineering[1, 2]. The increasing number of components in these systems leads to more frequent and complex interactions, greatly raising the probability of faults [3]. The complexity of the processes and the propagation of faults make fault classification increasingly difficult. The occurrence of faults can result in equipment downtime, production interruptions, and resource waste, leading to significant economic losses and even posing a threat to personnel safety [4]. Therefore, developing effective fault diagnosis methods is essential to ensure the safe and reliable operation of industrial systems.
With the deployment of a large number of sensors, industrial systems collect and store massive amounts of process data, and the exploration of data-driven fault diagnosis methods has become increasingly popular [5, 6]. Data-driven fault diagnosis methods include statistical methods and machine learning methods. The statistical methods for fault diagnosis [7] include principal component analysis (PCA), partial least squares (PLS), independent component analysis (ICA), Fisher discriminant analysis (FDA) [8, 9, 10, 11]. The machine learning methods for fault diagnosis include models such as support vector machine (SVM),k-nearest neighbors (KNN), and Bayesian network (BN) [12, 13, 14]. These methods have limitations when dealing with high-dimensional data and often require significant domain expertise for feature extraction. Compared to traditional machine learning methods, deep learning methods have advantages such as automatic feature extraction and higher accuracy [15]. Currently, deep learning methods applied to fault diagnosis in industrial systems include convolutional neural networks (CNN), stacked autoencoders (SAE), long-short-term memory networks (LSTM), temporal convolutional networks (TCN), and visual transformers (ViT). Wu et al. [16] constructed a fault diagnosis method based on deep convolutional neural networks (DCNN). This model incorporates convolutional layers, pooling layers, dropout and fully connected layers to extract feature representations. Liu et al. [17] integrated the sparse autoencoder with the denoising autoencoder to capture robust, sparse but intrinsic nonlinear features. Huang et al. proposed a fault diagnosis model named CNN-LSTM based on a combination of CNN and LSTM. The CNN layers and the LSTM layers are used to perform feature learning and capture time delay information respectively [18]. Zhou et al. proposed industrial process optimization vision transformer (IPO-ViT) fault diagnosis method based on ViT [19]. The self-attention mechanism in this model is used to perform global feature learning on the complex process signals. Zhu et al. combined the multi-scale feature extraction and the deep bidirectional mechanism to better extract the time series features [20]. These models primarily focus on fault diagnosis under a single operating mode. However, as the environmental conditions, loads, and production schedules change, industrial systems often switch between multiple operating modes. The significant distributional differences among monitoring data from multiple modes pose significant challenges for existing fault diagnosis methods.
Recently, some studies have focused on fault diagnosis in multimode processes. These works primarily address the challenge of distribution differences between the training set (source domain) and the testing set (target domain), which are referred to as cross-domain fault diagnosis. The techniques employed in cross-domain fault diagnosis methods are classified into data extension, learning strategy, and representation learning. Data extension promotes the learning of cross-domain generalizable feature representations by increasing the diversity of training data. The methods for increasing data diversity include data augmentation techniques (e.g., MixUp [21] and data transformations [22]) and data generation via deep neural networks [23, 24]. However, the effectiveness of these methods highly depends on the quality of the extended data. Representation learning mainly focuses on reducing the distribution discrepancy between different domains. These methods are categorized into adversarial-based methods and metric-based methods. In adversarial-based methods, adversarial learning is employed to deceive the discriminator, thereby capturing the domain-invariant features [25, 26, 27, 28, 29, 30]. Metric-based methods explicitly align features from different domains to capture generalizable feature representations. This is typically achieved by incorporating distance metrics into the loss function [31, 32, 33, 34, 35, 36]. Learning strategy exploits general learning methods (e.g., meta-learning [37]) to enhance the model’s generalization capability.
Although these fault diagnosis methods for multimode processes have achieved significant progress, there are still some open research gaps. On the one hand, current research focuses on fault diagnosis in multimode processes by capturing domain-invariant features. However, the varying amount of domain-invariant information contained in different time points is often neglected. Overemphasizing time points with limited domain-invariant information can degrade fault diagnosis performance. On the other hand, the operating mode of the samples is typically inaccessible, which makes methods relying on known operating modes inapplicable.
To address the above problems, a model named attention-based multi-scale temporal fusion network (AMTFN) is proposed. Specifically, the multi-scale depthwise convolution (MSDC) and gated recurrent unit (GRU) are employed to capture multi-scale contextual local features and long-short-term feature representations, respectively. Additionally, a temporal attention mechanism (TAM) is constructed to assign weights to deep features at different time steps, thereby enhancing the importance of specific time features that contain more domain-invariant information. The main contributions of this article are summarized as follows.
1) A significant scenario called uncertain-mode fault diagnosis is studied, where the operating mode of the samples is unavailable.
2) A fault diagnosis model named attention-based multi-scale temporal fusion network is proposed for uncertain-mode fault diagnosis.
3) Experiments on two datasets demonstrate that the AMTFN model can achieve effective fault diagnosis with a compact model size.
The rest of this paper is organized as follows. In Section 2, the problem description, related research areas and preliminary theoretical knowledge of depthwise convolution and GRU are introduced. The proposed method is explained in detail in Section 3. In Section 4, the experiments are conducted on the Tennessee Eastman (TE) process dataset and the three-phase flow dataset to validate the effectiveness of the proposed method. Finally, conclusions are drawn in Section 5.
2 Preliminaries
2.1 Problem formulation
Due to variations in environmental conditions or production schemes, industrial systems switch between multiple stable modes. Assume that a industrial system is designed to operate in $M$ modes, and the monitoring data collected from these modes is denoted as $D=\{(\bm{x}_{k}^{\text{o}},y_{k})\}_{k=1}^{N}\sim P_{XY}$, where $N$ is the number of samples for the $M$ modes. Here, $\bm{x}_{k}^{\text{o}}\in\mathbb{R}^{v}$ denotes the monitoring data at the $k$-th time point, with $v$ being the dimension of the measured variables, and $y_{k}\in\{1,...,L\}$ denotes the system health condition labels, including one normal category and $L-1$ fault categories. Although it is known that the samples are collected from $M$ modes, the specific mode to which each sample belongs is unknown. In the context of uncertain-mode fault diagnosis, it is assumed that fault data for the $M$ modes have been collected, with each mode containing samples for $L$ health conditions. The goal of uncertain-mode fault diagnosis is to build a model based on the monitoring data from these $M$ modes, such that the health condition of the system in these modes can be effectively identified.
2.2 Related research areas
There are several research fields related to uncertain-mode fault diagnosis (UMFD), including but not limited to: single mode fault diagnosis (SMFD), domain adaptation-based fault diagnosis (DAFD), and domain generalization-based fault diagnosis (DGFD). The differences between these methods and UMFD are summarized in Table 1.
Table 1: Comparison between UMFD and some related learning paradigms.
Settings Training set Testing set
Mode availability
of the $k$-th sample
SMFD $D^{i}$ $D^{i}$ $o_{k}=i$
DAFD $\{D^{i}\}_{i=1}^{S}\cup\{D^{’j}\}_{j=S+1}^{M}$$\{D^{j}\}_{j=S+1}^{M}$ $o_{k}\in\{1,…,M\}\text{ and }o_{k}=o_{0}$
DGFD $\{D^{i}\}_{i=1}^{S}$$\{D^{j}\}_{j=S+1}^{M}$ $o_{k}\in\{1,…,M\}\text{ and }o_{k}=o_{0}$
UMFD $\{D^{i}\}_{i=1}^{M}$$\{D^{i}\}_{i=1}^{M}$ $o_{k}\in\{1,…,M\}$
    $D’^{j}$ denotes the monitoring data without system health condition label. $D’^{j}=\{(\bm{x}_{k}^{j})\}_{k=1}^{N_{j}}$.
In SMFD, both the training and testing datasets are collected from the same mode. UMFD differs from SMFD in two main aspects. First, in UMFD, the operating mode of the sample is uncertain. This means that although it is known that the sample belongs to one of the $M$ operating modes, the specific mode is not identifiable. In the case where the operating modes of the sample are known, a fault diagnosis model can be built for each mode, and the corresponding model can be selected based on the sample’s operating mode. Second, UMFD involves monitoring data from multiple modes in both the training and test datasets. Taking the normal condition and fault 13 of the Tennessee Eastman (TE) process as an example, Fig. 1 presents the data distribution of two monitoring variables under six operational modes. Samples from the same health condition category may present complex data distributions due to mode changes, which puts higher requirements on the feature extraction capability of the fault diagnosis model.
Fig. 1.: Data distribution of monitoring variables under six modes in the TE process.
Cross-domain fault diagnosis aims to address the challenge posed by distribution discrepancies between the training set (source domain) and the testing set (target domain). It can be categorized into domain adaptation- based fault diagnosis (DAFD) and domain generalization-based fault diagnosis (DGFD) based on whether the target domain is accessible during the model training phase. Compared to cross-domain fault diagnosis, UMFD does not involve domain shift between the training and testing sets. However, the operating mode of the samples is inaccessible. The extraction of domain-invariant features is key to cross-domain fault diagnosis. UMFD aims to develop models that can effectively extract shared feature representations related to system health conditions from monitoring data across multiple modes. Therefore, the research on UMFD plays an important role in cross domain fault diagnosis.
2.3 Depthwise convolution
In traditional convolution, a convolution kernel performs operations across all input channels to generate a single output channel. In contrast, depthwise convolution assigns a single convolution kernel to each input channel and processes each channel independently, which greatly reduces computational complexity [38]. A depthwise convolution layer takes a feature map $\bm{x}\in\mathbb{R}^{w\times v}$ as input and produces a feature map $\bm{y}\in\mathbb{R}^{w\times v}$. $\bm{y}$ is formulated as $\bm{y}_{m,n}=\sum_{i}k_{m,i}\bm{x}_{m,n+i-1}$, where $k$ is the depthwise convolution kernel. The computation process of depthwise convolution is illustrated in Fig. 2.
Fig. 2.: Computation process of depthwise convolution.
2.4 GRU
Compared to the convolution neural network (CNN), the gated recurrent unit (GRU) have an advantage in processing time-series data. The GRU has fewer parameters and achieve faster convergence [39], and its structure is shown in Fig. 3.
Fig. 3.: Internal structure of GRU cell unit.
The GRU employs two gates, the reset gate and the update gate, to retain important information and discard irrelevant information. The output of GRU is formulated as follows [40],
$\bm{r}_{t}$ $=\sigma\left(\bm{W}_{r}\bm{b}_{t}+\bm{U}_{r}\bm{h}_{t-1}\right),$ (1)
$\tilde{\bm{h}}_{t}$ $=\tanh\left(\bm{W}\bm{b}_{t}+\bm{U}\left(\bm{r}_{t}\odot\bm{h}_{t-1}\right)\right),$
$\bm{z}_{t}$ $=\sigma\left(\bm{W}_{z}\bm{b}_{t}+\bm{U}_{z}\bm{h}_{t-1}\right),$
$\bm{h}_{t}$ $=\left(1-\bm{z}_{t}\right)\odot\bm{h}_{t-1}+\bm{z}_{t}\odot\tilde{\bm{h}}_{t},$
where $\sigma$ denotes the Sigmoid activation function. $\bm{W}$ and $\bm{U}$ are learnable weight matrices, which are shared across different time steps.
3 Proposed method
3.1 Overall structure of AMTFN
The overall structure of AMTFN is shown in Fig. 4. The model consists of three main modules: the feature extractor $E$, the feature fusion module $F$, and the classifier $C$. $E$ is used to capture deep feature representations, after which $F$ performs weighted fusion of features across different time steps. Finally, $C$ is employed to determine the health condition of the system.
Fig. 4.: Overall structure of AMTFN.
3.2 Feature extractor
The feature extractor consists of MSDC, channel-wise concatenation and GRU. MSDC is used to extract multi-scale contextual local features in parallel using convolution kernels of different sizes. Channel-wise concatenation is then employed to splice the multi-scale local features along the channel dimension, followed by GRU to further extract deep long-term and short-term feature representations.
To preserve dynamic temporal information, the samples from the previous $w-1$ time steps are extended to the current time step $\bm{x}_{k}=[\bm{x}_{k-w+1}^{\text{o}},\bm{x}_{k-w+2}^{\text{o}},...,\bm{x}_{k}^{\text{o}}]\in\mathbb{R}^{v\times w}$. To simplify the representation, let $\bm{x}_{k(i)}=\bm{x}_{k-w+i}^{\mathrm{o}}$, where $i=1,2,\ldots,w$. Then $\bm{x}_{k}=\left[\bm{x}_{k(1)},\bm{x}_{k(2)},\ldots,\bm{x}_{k(w)}\right]=\left[\bm{x}_{k-w+1}^{\mathrm{o}},\bm{x}_{k-w+2}^{\mathrm{o}},\ldots,\bm{x}_{k}^{\mathrm{o}}\right]$. $\bm{x}_{k}$ is fed into MSDC and then passed through channel-wise concatenation. The output of $\bm{x}_{k}$ after the above processing is formulated as,
$\bm{b}_{k}=\text{Concat}[\text{DC}_{1\times n}(\bm{x}_{k})],\quad n\in\{3,5,7,9\},$ (2)
where $\bm{b}_{k}\in\mathbb{R}^{4v\times w}$, DC denotes the depthwise convolution operation, and the subscript indicates the kernel size. In depthwise convolution, instance normalization is applied to weaken the mode features, and the ReLU function is used as the activation function. Concat represents concatenation along the channel dimension.
Then, the multi-scale features $\bm{b}_{k}$ are fed into the GRU, and the hidden state at the $t$-th time step is formulated as follows,
$\bm{h}_{k(t)}=\text{GRU}(\bm{b}_{k(t)},\bm{h}_{k(t-1)}).$ (3)
The output of the feature extractor for the $k$-th sample is formulated as, $\bm{h}_{k}=E\left(\bm{x}_{k}\right)=[\bm{h}_{k(1)},\bm{h}_{k(2)},...,\bm{h}_{k(w)}]$.
3.3 Feature fusion module
Considering that the amount of domain-invariant information varies across different time points, a temporal attention mechanism (TAM) is incorporated into the feature fusion module to enable the model to focus on features at critical time points. The feature fusion module takes the feature map $\bm{h}_{k}$ as input and infers a temporal attention map $\bm{a}_{k}$ through the TAM, as shown in Fig. 4. The temporal attention map is then used as a weight representing the importance of each time step to obtain the feature fusion result $\bm{f}_{k}$. The overall process of feature fusion can be summarized as follows,
$\bm{f}_{k}=F(\bm{h}_{k})=\sum_{t=1}^{w}\bm{g}_{k(t)}=\sum_{t=1}^{w}\bm{h}_{k(t)}\otimes\bm{a}_{k(t)},$ (4)
where $\otimes$ denotes element-wise multiplication.
The structure of the temporal attention mechanism is shown in Fig. 5. The relationships between different time steps of the features are utilized to generate the temporal attention map. Average pooling and standard deviation are used to aggregate information along the variable dimension of the feature map. These aggregated features are then fed into their respective fully connected layers to generate the average-pooling attention map $\bm{p}_{1}$ and the standard deviation attention map $\bm{p}_{2}$. The calculations for $\bm{p}_{1}$ and $\bm{p}_{2}$ are formulated as,
$\bm{p}_{1}$ $=F_{c2}\left(\sigma_{1}\left(F_{c1}(\text{AvgPool}\left(\bm{h}_{k}\right))\right)\right),$ (5)
$\bm{p}_{2}$ $=F_{c4}\left(\sigma_{1}\left(F_{c3}(\text{Std}\left(\bm{h}_{k}\right))\right)\right),$
where $\sigma_{1}$ denotes the ReLU function, AvgPool denote the average-pooling operations, Std denotes the calculation of standard deviation, $F_{c1}$, $F_{c2}$, $F_{c3}$ and $F_{c4}$ denote the fully connected layer.
Fig. 5.: Structure of TAM.
The average-pooling and standard deviation attention maps describe the attention to the feature map at different time steps from the perspectives of mean and variability, respectively. A convolution neural network is employed to fuse these two dimensions of attention and generate the final temporal attention. The temporal attention map $\bm{a}_{k}$ is formulated as,
$\bm{a}_{k}=\sigma_{2}\left(\text{Conv}\left(\text{Concat}\left(\bm{p}_{1};\bm{p}_{2}\right)\right)\right),$ (6)
where Conv denotes the convolution operation, $\sigma_{2}$ denotes ReLU function.
3.4 Model implementation and application process
The classifier takes the fused feature $\bm{f}_{k}$ as input and generates the final output through a fully connected layer followed by a softmax layer. The final output $\bm{c}_{k}$ is formulated as,
$\bm{o}_{k}$ $=F_{c}\left(\bm{f}_{k}\right)=\left[o_{k,1},o_{k,2},\ldots,o_{k,L}\right],$ (7)
$c_{k,l}$ $=C\left(f_{k,l}\right)=\operatorname{Softmax}\left(o_{k,l}\right)=\frac{\exp\left(o_{k,l}\right)}{\sum_{j=1}^{L}\exp\left(o_{k,j}\right)},$
where $l=1,2,...,L$, $\bm{o}_{k}$ denotes the output of the fully connected layer, $F_{c}$ represents the fully connected layer. $c_{k,l}$ represents the $l$-th dimension of the output for the $k$-th sample, which indicates the probability that the sample belongs to the $l$-th category. In addition, dropout is added before the fully connected layer to prevent overfitting. The cross-entropy loss function is used as the objective function for the fault diagnosis model, and it is formulated as,
$L=-\sum_{k=1}^{N_{b}}\sum_{l=1}^{L}y_{k,l}\log(c_{k,l})$ (8)
where $N_{b}$ denotes the number of samples in a batch, and $y_{k,l}$ denotes the $l$-th dimension of the one-hot encoding of the true class label for the $k$-th sample.
The process of applying the AMTFN model to fault diagnosis is illustrated in Fig. 6. It mainly consists of two steps: the first is data collection and pre-processing, and the second is model construction, optimization, and evaluation. The details are as follows:
(1) Data collection and pre-processing mainly includes data acquisition, labeling, standardization, sliding window processing, and data splitting. Firstly, the monitoring data across multiple modes are collected, and each sample is assigned a corresponding category label. Then, z-score normalization is applied to standardize samples across all modes. The standardized result is formulated as $(\bm{x}_{k}^{\text{o}}-\bm{\mu})/\bm{\sigma}$, where $\bm{\mu}$ and $\bm{\sigma}$ denote the mean and standard deviation of the normal samples across all modes. And sliding window processing is employed to preserve dynamic temporal information. Finally, the data are divided into training, validation, and testing sets.
(2) Model construction and optimization involves training the fault diagnosis model using the training set and selecting the optimal model based on the validation set. Model evaluation is performed on the test set to assess the overall performance of the fault diagnosis model.
Fig. 6.: The AMTFN model for uncertain-mode fault diagnosis.
4 Experiments
4.1 Evaluation metrics
Micro-F1, Macro-F1, Fault Diagnosis Rate(FDR), and False Positive Rate(FPR) are selected to evaluate the performance of the fault diagnosis model. To facilitate the presentation of the calculation formula for each evaluation metric, the confusion matrix for the $l$-th class is provided in Table 2, where each element represents the number of corresponding samples.
Table 2: Confusion matrix for the $l$-th class.
Predicted class is $l$
Predicted class is not $l$
Actual class is $l$ $TP_{l}$ $FN_{l}$
Actual class is not $l$ $FP_{l}$ $FN_{l}$
The Micro-F1 is computed as follows,
[…]
Micro-F1 is calculated at the dataset level and each sample has the equal weight.
The Macro-F1 is computed as follows,
[…]
Macro-F1 is calculated at the class level and each class has the equal weight.
The FDR and FPR are computed as follows,
[…]
$\text{FPR}_{l}$ $=\frac{\mathrm{FP}_{l}}{\mathrm{FP}_{l}+\mathrm{TN}_{l}}$
FDR, also known as Recall, measures the proportion of samples with the true label $l$ that are correctly classified as $l$. FPR measures the proportion of samples whose true label is not $l$ that are misclassified as $l$.
In addition, to evaluate the scalability of the fault diagnosis model, the number of parameters, training time, and testing time for each model are also used for comparison.
4.2 Implementation details
The detailed structural parameters of each module in AMTFN is shown in Table 3. In the experiments, the learning rate is set to 0.01. The epoch and batch size are set to 30 and 512, respectively.
Table 3: Structure of the modules in AMTFN.
Modules Symbols Type Input, output, kernel size
Feature $\text{DC}_{1\times 3}$ Depthwise conv ($v$$\times$64,$v$$\times$64,(3$\times$1)$\times$$v$)
extractor $E$ $\text{DC}_{1\times 5}$ Depthwise conv ($v$$\times$64,$v$$\times$64,(5$\times$1)$\times$$v$)
$\text{DC}_{1\times 7}$ Depthwise conv ($v$$\times$64,$v$$\times$64,(7$\times$1)$\times$$v$)
$\text{DC}_{1\times 9}$ Depthwise conv ($v$$\times$64,$v$$\times$64,(9$\times$1)$\times$$v$)
- GRU ($4v$$\times$64, 100$\times$64, -)
Feature $F_{c1}$ Fully connected (64, 64/$r$, -)
fusion $F$ $F_{c2}$ Fully connected (64/$r$, 64, -)
Conv Conv1d (3$\times$64, 1$\times$64, 3)
Classifier $C$ $F_{c3}$ Fully connected (100,$c$, -)
To validate the effectiveness of the proposed AMTFN, several advanced fault diagnosis models are applied for comparison. The models used for comparison are as follows:
(1) DCNN [16]. The DCNN model is constructed using convolutional layers, pooling layers, dropout, and fully connected layers, aiming to capture spatiotemporal features.
(2) CNN-LSTM [18]. In the CNN-LSTM model, the CNN layer is used for feature learning, while the LSTM layer is responsible for capturing time delay information.
(3) IPO-ViT [19]. In the IPO-ViT model, the Transformer architecture is employed for global feature learning of complex process signals.
(4) Multi-scale generative and adversarial Metric networks (MGAMN) [34]. MGAMN was developed for cross-domain fault diagnosis. In this model, adversarial training and metric learning strategy is designed to learn common feature representations.
Methods 1-3 have demonstrated superior performance in single-mode fault diagnosis, while Method 4 represents a state-of-the-art domain generalization technique. Advanced single-mode and domain generalization fault diagnosis methods are applied to uncertain mode fault diagnosis scenario to validate the challenges in such settings. The comprehensive comparison of all methods is conducted to evaluate the performance of the proposed model.
The model settings for the ablation experiments are shown in Table 4.
Table 4: Model settings for the ablation experiments.
MSDC GRU TAM
A1 $\checkmark$
A2 $\checkmark$
A3 $\checkmark$ $\checkmark$
A4 $\checkmark$ $\checkmark$
A5 $\checkmark$ $\checkmark$
4.3 Case one: TE process
4.3.1 Dataset description and fault diagnosis tasks
The TE process simulation model was developed based on a real chemical process, which has become a benchmark for fault diagnosis and process monitoring [41]. It mainly consists of five components: the reactor, condenser, vapor-liquid separator, recycle compressor and product stripper. Bathelt et al. [42] introduced additional process measurements and disturbances to the TE process, which were implemented using SIMULINK. The overall structure of the extended TE process is shown in Fig. 7. The TE process includes 41 measured variables and 12 manipulated variables. It is capable of simulating 28 faults under 6 modes. Liu et al. obtained the multimode fault diagnosis datasets by adjusting the parameters of the TE process simulation models [43]. Since fault 3 and fault 16 are analogous with the normal operating conditions [44], they are not included in this study. The system health condition categories, modes and task settings are presented in Table 5 and Table 6. In tasks T1 to T6, each task’s dataset consists of data from five modes, with the corresponding positions marked as $\checkmark$.
Fig. 7.: P&ID of the revised process model [42].
Table 5: Faults of TE process used in uncertain-mode fault diagnosis [45].
No. Description
N Normal
F1 A/C feed ratio, B composition constant (stream 4)
F2 B composition, A/C ratio constant (Stream 4)
F4 Reactor cooling water inlet temperature
F5 Condenser cooling water inlet temperature
F6 A feed loss (stream 1)
F7 C header pressure loss - reduced availability (stream 4)
F8 A, B, C feed composition (stream 4)
F9 D feed temperature (stream 2)
F10 C feed temperature (stream 4)
F11 Reactor cooling water inlet temperature
F12 Condenser cooling water inlet temperature
F13 Reaction kinetics
F14 Reactor cooling water valve
F15 Condenser cooling water valve
F17-20 Unknown
F21 The valve fixed at steady state position
Table 6: Modes of TE process used in uncertain-mode fault diagnosis.
No. G/H mass ratio and production rate T1 T2 T3 T4 T5 T6
M1 50/50, G: 7038kg/h, H: 7038kg/h $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
M2 10/90, G: 1408kg/h, H: 12669kg/h $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
M3 90/10, G: 10000kg/h, H: $1$kg/h $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
M4 50/50, maximum production rate $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
M5 10/90, maximum production rate $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
M6 90/10, maximum production rate $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
The TE dataset used in the experiment was generated through SIMULINK simulation of the TE process, covering 19 system health state categories across 6 modes. The simulation time was set to 100 hours, the sampling interval was set to 3 minutes and and the fault injection time was set to 30 hours. Note that during fault simulations in some modes, the simulation process might stop abruptly, resulting in fewer samples for some categories, leading to class imbalance. In sliding window processing, the window size $w$ was set to 64. For the data from each mode, a hierarchical splitting method was applied to divide the samples into training, validation, and testing sets in an 8:1:1 ratio.
4.3.2 Experimental results and analysis
The Micro F1 and Macro F1 scores of different methods on the TE process dataset are shown in Table 7 and Table 8, respectively. The proposed AMTFN achieves the highest average Micro F1 and average Macro F1 scores among the five fault diagnosis models, with values of 0.9792 and 0.9803, respectively. The DCNN has the lowest average Micro F1 and average Macro F1 scores, indicating that relying solely on convolution operations struggles to capture the complex features of multimode data distributions. Compared to the CNN-LSTM, IPO-ViT, and MGAMN models, AMTFN achieves a 3.18%, 9.08%, and 14.19% improvement in average Micro F1, respectively, and a 2.98%, 8.48%, and 15.47% improvement in average Macro F1, respectively. Taking Task T1 as an example, the two models with the highest average Micro F1 scores among the comparative models are selected for further comparison. The FDR and FPR scores of different methods on the TE process dataset are shown in Table 9. The proposed AMTFN model achieved the highest FDR, with FDR values exceeding 0.9 for both the N and F21 categories, whereas the comparative models showed FDR values below 0.8 for these categories. In addition, the AMTFN model achieved the lowest FPR. This validates that the proposed AMTFN model can effectively capture deep features in multimode distributed data, enabling efficient fault diagnosis under uncertain modes.
Table 7: Micro F1 scores of different methods on the TE process dataset.
Task Model
DCNN CNN-LSTM IPO-ViT MGAMN AMTFN
T1 0.7187 0.9595 0.8941 0.8718 0.9824
T2 0.6735 0.9236 0.9015 0.8501 0.9758
T3 0.6292 0.9526 0.8939 0.8376 0.9789
T4 0.6786 0.9709 0.9176 0.8602 0.9814
T5 0.6088 0.9488 0.8989 0.8326 0.9777
T6 0.6258 0.9385 0.8802 0.8356 0.9789
Avg 0.6558 0.9490 0.8977 0.8480 0.9792
Table 8: Macro F1 scores of different methods on the TE process dataset.
Task Model
DCNN CNN-LSTM IPO-ViT MGAMN AMTFN
T1 0.6159 0.9622 0.9008 0.8809 0.9835
T2 0.5795 0.928 0.9077 0.8621 0.9774
T3 0.5305 0.955 0.8992 0.8503 0.9795
T4 0.5777 0.9723 0.9216 0.8663 0.9823
T5 0.5339 0.9518 0.9055 0.8422 0.9791
T6 0.4996 0.9424 0.8876 0.8492 0.9802
Avg 0.5562 0.9520 0.9037 0.8585 0.9803
Table 9: FDR and FPR scores of different methods on the TE process dataset.
CNN-LSTM IPO-ViT AMTFN
Class FDR FPR FDR FPR FDR FPR
N 0.7686 0.0136 0.5443 0.0281 0.9186 0.0061
F1 0.9983 0.0002 0.9948 0.0009 1 0.0002
F2 0.9986 0 0.9914 0.0004 0.9986 0.0004
F4 0.9986 0 0.9957 0 1 0
F5 0.9957 0 0.9929 0.0003 0.9971 0
F6 1 0 1 0 1 0
F7 1 0 1 0 1 0
F8 0.9914 0.0004 0.9671 0.0007 0.9929 0.0003
F9 0.93 0.0054 0.7429 0.0176 0.9829 0.0013
F10 0.98 0.0019 0.8929 0.0085 0.9886 0.0008
F11 0.9857 0.0002 0.9743 0.0002 0.9957 0.0002
F12 0.9957 0.0003 0.99 0.0004 0.9971 0.0002
F13 0.9814 0.0008 0.9686 0.002 0.9857 0.0013
F14 0.9929 0.0003 0.98 0.0019 0.9943 0
F15 0.9914 0.001 0.6943 0.0187 1 0.0001
F17 0.9757 0.0008 0.9771 0.001 0.9771 0.0002
F18 0.95 0.0019 0.9357 0.001 0.9457 0.0014
F19 0.98 0.001 0.9586 0.0018 0.9957 0.0001
F20 0.9771 0.0016 0.9757 0.0015 0.9857 0.0015
F21 0.7514 0.0136 0.4414 0.0269 0.9143 0.0045
AVG 0.9621 0.0022 0.9009 0.0056 0.9835 0.0009
T-distributed stochastic neighbor embedding (t-SNE) is used to visualize the feature distributions learned by each model. Taking task T1 as an example, the visualization results are shown in Fig. 8. From Fig. 8(a) and (d), it can be seen that the features learned by CNN and MGAMN reveal poorly separated clusters. Fault 14 is selected to compare CNN-LSTM, IPO-ViT, and the proposed model. The samples belonging to Fault 14 are highlighted with dashed boxes in Fig. 8(b), (c), and (e). It is evident that the features extracted by CNN-LSTM and IPO-ViT are too scattered, whereas AMTFN effectively clusters the samples belonging to this category. This demonstrates that the features extracted by AMTFN exhibit minimal differences across modes, highlighting its superior cross-mode performance.
Fig. 8.: Feature map visualization results via t-SNE on task T1.
To evaluate the scalability of different models, the number of parameters, training time per epoch, and testing time per epoch for each model are presented in Table 10. AMTFN has the smallest number of parameters and the shortest training time, with testing time exceeding that of MGACN by only 0.28 seconds. This is because the GRU in AMTFN limits parallel processing, which adds the computation time. However, AMTFN demonstrates unparalleled advantages in inferring system health conditions. The computational complexity of DCNN, CNN-LSTM, and IPO-ViT is significantly higher than that of AMTFN, which highlights the advantages of AMTFN in practical applications.
Table 10: Parameter number, training time and test time of different methods on the TE process dataset.
Metrics Model
DCNN CNN-LSTM IPO-ViT MGAMN AMTFN
Parameter
number (M)
11.41 0.91 26.39 0.60 0.48
Training
time (s/epoch)
18.73 11.17 86.26 6.19 4.87
Test time
(s/epoch)
2.76 2.59 5.92 2.11 2.39
4.3.3 Ablation study
To verify the effectiveness of the MSDC, GRU, and TAM, ablation experiments were conducted. The Micro F1 scores of different methods are shown in Table 11. The average Micro F1 score of A2 is 17.56% higher than that of A1, indicating that the GRU is superior to the CNN in temporal feature extraction. The model A3 and A4 are obtained by adding TAM to A1 and A2, respectively (see Table 4). The average Micro F1 score of A3 is 10.19% higher than that of A1, and A4 is 1.63% higher than that of A2. This demonstrates that TAM effectively fuses features extracted by both CNN and GRU. A5 combines A1 and A2, meaning that both the MSDC and GRU are used for feature extraction. The average Micro F1 score of A5 is 20.35% higher than that of A1 and 2.38% higher than that of A2. This demonstrates that combining MSDC and GRU enhances the ability to capture the complex structures within the data. AMTFN is equivalent to adding GRU to A3, or adding MSDC to A4, or adding TAM to A5. The average Micro F1 score of TAGRUN is 9.35% higher than A3, 0.84% higher than A4, and 0.11% higher than A5. This indicates that GRU, MSDC, and TAM complement each other, collectively improving the model’s performance in uncertain-mode fault diagnosis.
Since the improvement in AMTFN’s average Micro F1 score over A4 and A5 is relatively small, the feature maps extracted from A4 and A5 are visualized using t-SNE. Taking Task T1 as an example, the visualization results are shown in Fig. 9. Similarly, the samples belonging to Fault 14 are highlighted with dashed circles. Although A4 achieves a Micro F1 score of 0.9707, the features of samples from different categories exhibit poor separability. A5 effectively clusters the features of samples from different categories into different groups. However, the features of samples within the same category exhibit significant dispersion. This may be because A5 learns mode-specific features, causing each operation mode and system health condition category pair to be clustered into a unique group as tightly as possible. For uncertain-mode fault diagnosis, it is essential to ensure both the separability of samples from different categories and the compactness of samples from the same category. The proposed AMTFN model achieves better intra-category compactness and maintains a reasonable degree of inter-category separability. This demonstrates that focusing on critical time points through TAM may help the model capture invariant features across different modes, thereby reducing feature differences among samples of the same category.
Table 11: Micro F1 scores of different methods on the TE process dataset.
Task Model
A1 A2 A3 A4 A5 AMTFN
T1 0.8118 0.9466 0.9003 0.9707 0.9793 0.9824
T2 0.8259 0.9595 0.8889 0.9743 0.9748 0.9758
T3 0.8058 0.9517 0.8691 0.9699 0.9782 0.9789
T4 0.8008 0.9612 0.9259 0.9743 0.9808 0.9814
T5 0.8292 0.9599 0.9233 0.9692 0.9765 0.9777
T6 0.8027 0.9533 0.8657 0.9676 0.9790 0.9789
Avg 0.8127 0.9554 0.8955 0.9710 0.9781 0.9792
Fig. 9.: Feature map visualization results via t-SNE on task T1.
4.4 Case two: three-phase flow facility
4.4.1 Dataset description and fault diagnosis tasks
In the field of fault diagnosis, the dataset from three-phase flow facility [46] at Cranfield University has been used extensively as a benchmark for evaluating the performance of various methods. The three-phase flow facility is designed to provide controlled flows of water, oil, and air. Its structure is shown in Fig. 10. This facility captured 24 process variables and simulated six faults. The operation mode settings and the dataset used in the experiments are presented in Table 12, where different water flow rates and air flow rates correspond to different operating modes. The partitioning ratios of the training, validation, and test sets, as well as the sliding window settings, are consistent with those set in the TE process.
Fig. 10.: Sketch of the three-phase flow facility. [46].
Table 12: Faults of three-phase flow facility used in uncertain-mode fault diagnosis.
Data set index Water flow rate (kg/s) Air flow rate ($\text{m}^{3}$/s)
1.2, 4.2 2 0.0417
1.3, 3.3, 4.3 3.5 0.0208
2.2, 3.2 2 0.0278
2.3 3.5 0.0417
No. Description Data set index
N Normal 1.2,1.3,2.2,2.3,3.2,3.3,4.2,4.3
F1 Air line blockage 1.2,1.3
F2 Water line blockage 2.2,2.3
F3 Top separator input blockage 3.2,3.3
F4 Open direct bypass 4.2,4.3
4.4.2 Experimental results and analysis
The FDR, FPR, Micro F1 and Macro F1 scores of different methods on the three-phase flow facility dataset are shown in Table 13. The proposed AMTFN achieves the highest FDR, average Micro F1, and average Macro F1 scores, all with values of 1.0. In addition, the AMTFN model achieved the lowest FPR, with a value of 0. Compared to DCNN, CNN-LSTM, IPO-ViT, and MGAMN, AMTFN’s average FDR is higher by 25.15%, 0.01%, 6.93%, and 8.51%, respectively; its average Micro F1 is higher by 12.21%, 0.03%, 3.75%, and 4.68%, respectively; and its Macro F1 is higher by 35.10%, 0.03%, 6.18%, and 7.68%, respectively. For the system health condition category N, the FDR scores of DCNN, IPO-ViT, and MGAMN are all below 0.72, whereas both AMTFN and CNN-LSTM achieve a score of 1. Additionally, CNN-LSTM achieves an FDR of 0.9993 for category F3, whereas AMTFN still achieves a score of 1. Overall, the AMTFN model outperforms the other models across all four metrics, further validating its superiority in uncertain-mode fault diagnosis.
Table 13: FDR, FPR, Micro F1 and Macro F1 scores of different methods on the three-phase flow facility dataset.
Metrics Class Model
DCNN CNN-LSTM IPO-ViT MGAMN AMTFN
FDR N 0 1 0.7135 0.6324 1
F1 1 1 1 1 1
F2 1 1 0.9653 0.9838 1
F3 0.9952 0.9993 0.9972 0.9917 1
F4 1 1 1 1 1
Avg 0.7990 0.9999 0.9352 0.9216 1
FPR N 0 0.0003 0.0061 0.0061 0
F1 0.0358 0 0.0284 0.0301 0
F2 0.0297 0 0.0007 0.0016 0
F3 0.0134 0 0.003 0.0134 0
F4 0.0551 0 0.0059 0.0063 0
Avg 0.0268 0.0001 0.0088 0.0115 0
[…]
The feature visualization results obtained using t-SNE are shown in Fig. 11. Although CNN-LSTM achieves a Micro F1 score of 0.9997, a comparison of Fig. 11(b) and (e) clearly shows that the features extracted by CNN-LSTM exhibit weaker intra-class compactness and inter-class separability compared to those extracted by AMTFN. From Fig. 11(a), (c) and (d), it can be seen that the features extracted by these models form poorly separable clusters. In contrast, AMTFN achieves promising results, with features of samples from the same class exhibiting regional distributions and clear inter-class distances.
Fig. 11.: Feature map visualization results via t-SNE.
The parameter number, training time per epoch, and testing time per epoch for each model are presented in Table 14. The experimental results are similar to those obtained in TE process. This demonstrates that AMTFN is sufficiently lightweight, proving its advantage in practical applications.
Table 14: Parameter number, training time and test time of different methods on the three-phase flow facility dataset.
Metrics Model
DCNN CNN-LSTM IPO-ViT MGAMN AMTFN
Parameter
number
4.42 0.69 26.35 0.58 0.63
Training
time (s/epoch)
3.87 3.81 22.1 2.97 3.21
Test time
(s/epoch)
2.12 2.1 3.11 1.91 2.06
5 Conclusion
This article proposed a model named AMTFN for uncertain-mode fault diagnosis. The AMTFN stands out by focusing on critical time points with higher cross-mode shared information, significantly enhancing its ability to extract domain-invariant features. This improvement leads to superior performance in uncertain-mode fault diagnosis tasks. The proposed method was evaluated on the TE process and the three-phase flow facility. The experiments demonstrate that AMTFN exhibits significant superiority in terms of fault diagnosis performance, visualization, and lightweight design.
However, uncertain-mode fault diagnosis has certain limitations. It assumes that there is no distribution shift between the training and testing datasets. The industrial systems may experience new faults or operate in new modes. Therefore, exploring open-set uncertain-mode fault diagnosis and addressing the generalization problem in uncertain-mode fault diagnosis could be necessary future directions.